OpenStudy (anonymous):

Natalia constructed a right triangle with two side lengths measuring 3 and 5 units. She constructed a second right triangle similar to the first one, with the two corresponding sides measuring 5.1 and x units in length, respectively. What is the missing length, x, of the second triangle The figures shown are similar. What is the scale factor.

OpenStudy (agl202):

What do u think?

OpenStudy (agl202):

The correct answer is 8.5!!!

OpenStudy (welshfella):

corresponding sides are in the same ratio therefore 3 / 5.1 = 5 / x x = (5*5.1) / 3